OpenStudy (anonymous):

how to solve cos (4x)=-1/2 ?

OpenStudy (anonymous):

find a number (angle if you like) whose cosine is \(-\frac{1}{2}\) which should not be too hard

OpenStudy (anonymous):

if it is not obvious, look at the unit circle on the last page of the attached cheat sheet see what angle has a first coordinate of \(-\frac{1}{2}\)

OpenStudy (anonymous):

wouldnt that be both pi/3 and 5pi/3??? would this yield two general answers?

OpenStudy (anonymous):

no \(\cos(\frac{\pi}{3})=\frac{1}{2}\) you want \(-\frac{1}{2}\)

OpenStudy (anonymous):

try \(\frac{2\pi}{3}\) and \(\frac{4\pi}{3}\)

OpenStudy (anonymous):

this tells you \[4x=\frac{2\pi}{3}\] and so \[x=\frac{\pi}{6}\]

OpenStudy (anonymous):

or \[4x=\frac{4\pi}{3}\] and therefore \[x=\frac{\pi}{3}\]

OpenStudy (anonymous):

oh ya! so then it still gives the answers of 2pi/3 +2pi n and 4pi/3 +2pi n ?

OpenStudy (anonymous):

wat about the other value of x?

OpenStudy (anonymous):

no you are solving for \(x\) not for \(4x\)

OpenStudy (anonymous):

gotta divide your answers by 4

OpenStudy (anonymous):

\[4x=\frac{2\pi}{3}+2n\pi\] \[x=\frac{\pi}{6}+\frac{1}{2}n\pi\]

OpenStudy (anonymous):

I mean like cos(2pi/3) and cos(4pi/3)...

OpenStudy (anonymous):

would u also plug in the 4pi/3, i mean.

OpenStudy (anonymous):

hold on lets go slow

OpenStudy (anonymous):

\[\cos (\frac{2\pi}{3})=-\frac{1}{2}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and your job was to solve \[\cos(4x)=-\frac{1}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this tells you that \[4x=\frac{2\pi}{3}+2n\pi\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

now divide by \(4\) to solve for \(x\) and get \[x=\frac{\pi}{6}+\frac{1}{2}n\pi\]

OpenStudy (anonymous):

or however you wish to write it

OpenStudy (anonymous):

So my question is,

OpenStudy (anonymous):

are there two answers, since there are two different angles that cosine can be if it's -1/2?

OpenStudy (anonymous):

right now repeat the process with \[4x=\frac{4\pi}{3}+2n\pi\]

OpenStudy (anonymous):

yes there are two answers, or rather two formulas for infinite number of answers

OpenStudy (anonymous):

pi/6 + pi n/2 and pi/3 + pi n/2 yes? :)

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

>o< ok then that clears things up~ and the +2pi n thing would be just +pi n for tangent and cotangent, right? I just need to clear that up...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is there a plus or minus thing involved in solving trig equations? I went in to tutoring this morning and heard something about having to put plus or minus in the process, but I had to go early, so I wasn't able to ask what scenario that was in... I think I heard something about cosine...

OpenStudy (anonymous):

hello?